In a dielectric material with ∈_r=3.6 , D= 285 nC/m^2. Find the magnitudes of E, P, and electric susceptibility


1
Expert's answer
2021-05-17T13:31:59-0400

As we know,


"D=\epsilon_0\epsilon_rE,\\\space\\
E=\frac{D}{\epsilon_0\epsilon_r}=8941\text{ V/m}."

On the other hand, we know that


"D=\epsilon_0E+P,\\
P=D-\epsilon_0E=205.8·10^{-9}\text{ C/m}^2."

Find electric susceptibility:


"P=\epsilon_0\chi_eE,\\\space\\
\chi_e=\frac{P}{\epsilon_0E}=2.6."